Find the uncertainty in position of electron when the uncertainty in its velocity is 0.0005 % of velocity of light.


1
Expert's answer
2020-12-08T05:31:57-0500

∆x∆p = h/4π

∆x (m∆v) = h/4π

∆x = h/(4πm∆v)

mass of electron is 9.1×10-31 kg

∆v is given = 0.0005% × 3× 108 m/s= 0.000005 × 3× 108 m/s = 15 × 103m/s

h = 6.626× 10-34

on substituting these values we get

∆x = 3.86 × 10-6 m